x + 2y - 7z = -1
3x + 7y - 24z = 6
x + 4y - 12z = 26

Answer by stanbon(75887)

x + 2y - 7z = -1
3x + 7y - 24z = 6
x + 4y - 12z = 26
-------------------------
I used the Matrix capability of a TI calculator to find:
x = -10
y = 36
z = 9
==================
Cheers,
Stan H.